OpenStudy (anonymous):

find the vertex. y=-x^2+2x+4

OpenStudy (tkhunny):

Complete the square! y = -(x^2 - 2x + ____) + 4 + ____

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

I would do -b/2a

OpenStudy (anonymous):

So, a=-1,b=2, c=4, now you are putting -b/2a -2/-2=1, and then you will plug the value of 1 to find the corresponding y coordinate

OpenStudy (anonymous):

\[x^2+ 2x+4\\ ( x^2+ 2x )+4\\ x^2+ 2x+( \frac{2}{2})^2 ) + 4 - ( \frac{2}{2})^2 ) \\ ( x + 1)^2 + 3 \]

OpenStudy (anonymous):

\[ \text{ So, here, the Vertex} = ( -1,3) \]

OpenStudy (anonymous):

luis its -x^2

OpenStudy (tkhunny):

If you Complete the Square, like Luis_Rivera just showed (except for the typo), you can PROVE -b/(2a)! Why would you NOT want to do that?

OpenStudy (anonymous):

Because I would rather do -b/2a than the long and complicated way.

OpenStudy (anonymous):

oh ok i got it thanks

OpenStudy (anonymous):

Your welcome :D

OpenStudy (anonymous):

lol